A small particle has charge -5.00 uC and mass



2.00 x 10-4 kg. It moves from point A, where the electric poten-



tial is VA = +200 V, to point B, where the electric potential is



VB = +800 V. The electric force is the only force acting on the



particle. The particle has speed 5.00 m/s at point A. What is its



speed at point B? Is it moving faster or slower at B than at A?

1
Expert's answer
2022-04-22T16:43:43-0400

"\frac{mv_B^2}2=q∆\varphi+\frac{mv_A^2}2,"

"v_B=\sqrt{\frac{2q∆\varphi}m+v_A^2}=8.7~\frac ms."